OpenStudy (anonymous):

when the area of an expanding square is increasing three times as fast as its side is increasing the side is

OpenStudy (blockcolder):

Let A be the area and s be the side of the square. From your given condition we have \[\frac{dA}{dt}=3\frac{ds}{dt}\] But we know that A=s^2 so \[\frac{dA}{dt}=2s \frac{ds}{dt}\] Chugging this back to the first equation, we have: \[2s\frac{ds}{dt}=3\frac{ds}{dt}\\ 2s=3\\ s=\frac{3}{2}\]